OpenStudy (anonymous):

Betsy is testing whether school is more enjoyable when students are making high grades. She asked 150 students if they enjoyed school and whether their GPA was above or below 3.5. She found that 35 of the 60 students with a GPA above 3.5 reported that they enjoyed school, and 15 of the 90 students with a GPA below 3.5 reported that they enjoyed school. What is the probability that a student with a GPA below 3.5 does not enjoy school?

OpenStudy (kropot72):

The probability that a student with a GPA below 3.5 enjoys school is 15/90 = 1/6. Therefore the probability that a student with a GPA below 3.5 does not enjoy school is given by: \[1 - \frac{1}{6}=\frac{5}{6}\ or\ 0.833\]